# FEUDA: Frustratingly Easy Prompt Based Unsupervised Domain Adaptation

### v1

@@ Abstract
A major thread of unsupervised domain adaptation (UDA) methods uses unlabeled data from both source and target domains to learn domain-invariant representations for adaptation. However, these methods showcase certain limitations, encouraging the use of self-supervised learning through continued pre-training. The necessity of continued pre-training or learning domain-invariant representations is still unclear in the prompt-based classification framework, where an input example is modified by a template and then fed into a language model (LM) to generate a label string. To examine this new paradigm of UDA in the prompt-based setup, we propose a frustratingly easy UDA method (FEUDA) that trains an autoregressive LM on both unlabeled and labeled examples using two different instruction-tuning tasks. Specifically, the first task trains the LM on unlabeled texts from both domains via masked language modeling (MLM), and the other uses supervised instruction-tuning on source-labeled data for classification. We conduct extensive experiments on 24 real-world domain pairs to show the effectiveness of our method over strong domain-invariant learning methods. Our analysis sheds light on why masked language modeling improves target-domain classification performance in prompt-based UDA. We discover that MLM helps the model learn both semantic and background knowledge of a domain, which are both beneficial for downstream classification.
@@ 1 Introduction
Despite recent advancements in the pre-training of language models, these models are still fragile under certain kinds of data distribution shifts, making their real-world applications challenging Ribeiro et al. (2020). The problem of unsupervised domain adaptation (UDA) aims to leverage learned knowledge from a labeled source domain to an unlabeled target domain Pan and Yang (2010); Ganin and Lempitsky (2015); Long et al. (2015).
A vast class of existing UDA methods attempts to learn representations that are invariant across domains Tzeng et al. (2014); Ganin et al. (2016); Wu and Shi (2022); Guo et al. (2022). The rationale is that when the learned representations from both domains cannot be distinguished by a classifier and the classifier performs well on the source domain, it will also exhibit strong performance on the target domain. However, previous work has shown that domain-invariance is insufficient for adaptation to the target domain Zhao et al. (2019), and is also prone to instability issues Han and Eisenstein (2019); Kashyap et al. (2021). This has encouraged the emergence of self-supervised approaches through language model pre-training. Variants of continued pre-training have proven to be effective and stable for adapting pre-trained LMs to labeled and unlabeled downstream tasks Gururangan et al. (2020); Karouzos et al. (2021).
However, the trade-offs between continued pre-training and learning domain-invariant representations for UDA are both unexplored in the prompt-based classification framework Gao et al. (2021); Liu et al. (2023). In such a framework, input examples are modified using instruction templates and then fed into a language model (LM) to generate the label text based on the constructed instruction. This process bears resemblance to instruction tuning Wei et al. (2022), except that the training is done on a single task, and that the end goal is adaptation to a specific task, rather than generalization to unseen tasks. In this paper, we call this new paradigm prompt-based UDA and examine two research questions: Can we utilize unlabeled data to construct useful instruction-tuning tasks for UDA? Is domain invariance still necessary in this paradigm?
To answer these questions, we present a “frustratingly easy” UDA method (termed FEUDA), which smooths out the transition between pre-training and adaptation by two instruction-tuning tasks using prompt templates. First, unlabeled texts from both source and target domains are modified by a prompt template and then used to train an autoregressive LM to perform the masked language modeling (MLM) task. Next, the LM is instruction-tuned on labeled source texts using another template for the classification task.
Through extensive experiments on 40 real-world domain pairs, various adaptation methods, and few-shot learning setups, we show that FEUDA is competitive for UDA and even outperforms methods that explicitly promote domain invariance (Section 6.1). Additionally, our analysis sheds light on how masked language modeling improves classification performance on the unlabeled target domain (Section 6.2). We discover that MLM helps the model learn both semantic and background knowledge of a domain, which are both beneficial for downstream classification. Our main contributions can be summarized as follows:
  1. 1.
    We introduce prompt-based UDA, a new UDA setting where the discriminative prediction is converted into a generative task, enabling multi-task adaptation as well as the reuse of all language model parameters. We empirically analyze continued pre-training and domain invariance based UDA methods in this setting.
  2. 2.
    We propose FEUDA, a simple and effective UDA approach for prompt-based classification. Through extensive experiments, we show that FEUDA is competitive and outperforms the domain-invariant learning approach. We establish the generalizability of FEUDA across various models, adaption methods and limited data settings, confirming that our approach remains powerful in these settings.
  3. 3.
    We conduct an analysis understanding the impact of the MLM task in a UDA setup and discover that MLM helps the model learn both semantic and background knowledge of a domain, both of which are beneficial for downstream classification.
2 Related Work
Ramponi and Plank (2020) categorize UDA methods into two general classes: Model-centric and Data-centric methods. This work focuses on a new data-centric UDA method in a prompt-based setup.
Model-centric UDA Methods
This line of study involves augmenting the feature space (Blitzer et al., 2006; Pan et al., 2010; Ziser and Reichart, 2018, inter alia), editing models though weight interpolation Matena and Raffel (2022); Cai et al. (2023); Wortsman et al. (2022); Ilharco et al. (2022), or altering the loss function and model architecture. One typical framework aims to minimize $\mathcal{H}\Delta\mathcal{H}$ divergence Ben-David et al. (2010) between the source and target domain features, through adversarial training (Tzeng et al., 2014; Ganin et al., 2016; Tzeng et al., 2017; Wu and Shi, 2022; Guo et al., 2022, inter alia) or through minimizing measures of domain similarity Bousmalis et al. (2016); Ge et al. (2023); Malik et al. (2023). However, past work has shown that domain-invariance is weak constraint for adaptation to the target domain Zhao et al. (2019); Karouzos et al. (2021), could introduce domain-specific hyperparameters Trung et al. (2022), and is also prone to instability issues Han and Eisenstein (2019); Sun et al. (2019); Wilson and Cook (2020); Kashyap et al. (2021).
Data-centric UDA Methods
The limitations of invariance-based model-centric methods have encouraged the emergence of alternate approaches, largely based on self-supervised learning through contrastive learning Kumar et al. (2022); Shen et al. (2022); Long et al. (2022), pseudo-labeling (Zhou and Li, 2005; Ruder and Plank, 2017, inter alia) or language model pre-training. Despite not being directly useful to certain downstream tasks Uppaal et al. (2023), Masked Language Modelling (MLM) has been used for adaptation to labeled tasks, in both full fine-tuning Gururangan et al. (2020); Lee et al. (2020); Gao et al. (2021) and PEFT setups Kim et al. (2021); Hung et al. (2023). A smaller body of work has explored the utility of MLM in a UDA setup Han and Eisenstein (2019); Zhang et al. (2021b); Karouzos et al. (2021). This class of methods is more stable than invariance-based methods, but often requires additional compute for extended pre-training.
Prompt-based UDA
The emergence of large language models (Brown et al., 2020; Scao et al., 2022; Touvron et al., 2023, inter alia) introduced the concept of instruction tuning, where a language model is trained on strings of input-output pairs, often using instruction-specific templates Zhang et al. (2023). Inspired by this, we introduce prompt-based UDA, i.e., a new paradigm of data-centric UDA using a prompt-based classifier that casts the discriminative classification task into a generative next-token prediction task. This prompt-based UDA formulation provides two unique benefits compared to traditional UDA approaches: 1) the prompt-based classifier can reuse all model parameters for adaptation, without requiring any task-specific architectural changes; 2) furthermore, this also enables multi-task instruction tuning bridging the gap between pre-training and adaptation, as the pre-training and fine-tuning phases of the exact model can be naturally coupled together by using different instruction prompts.
Our study aims to extend the existing body of data-centric UDA methods by examining the behavior of multi-task instruction tuning for adaptation and the necessity of learning domain invariant representations in this new UDA paradigm.
3 Preliminaries: The UDA Problem
@@
We consider a text classification task, where $\mathcal{X}$ is the input space of all text sentences and $\mathcal{Y}=\{1,...K\}$ is the label space. In the UDA problem, we have access to a source labeled dataset $\mathcal{D}_{\text{src}}=\{(x_{i},y_{i})\}_{i=1}^{N}$ consisting of samples from a joint distribution $P_{\mathcal{XY}}^{\text{(src)}}$, and a target unlabeled dataset $\mathcal{D}_{\text{tgt}}=\{x_{j}\}_{j=1}^{M}$ sampling from a target input distribution $P_{\mathcal{X}}^{\text{(tgt)}}$. We further denote $P_{\mathcal{X}}^{\text{(src)}}$ as the marginal distribution of $P_{\mathcal{XY}}^{\text{(src)}}$ on $\mathcal{X}$, where $P_{\mathcal{X}}^{\text{(src)}}\neq P_{\mathcal{X}}^{\text{(tgt)}}$. The goal of UDA is to learn a function $f:\mathcal{X}\to\mathcal{Y}$ such that the error rate $\mathbb{E}_{x\sim P_{\mathcal{X}}^{\text{(tgt)}}}\;\mathbbm{1}[f(x)\neq y]$ is minimized.
4 FEUDA Method
@@
In this section, we introduce our framework FEUDA, a simple and effective two-phase training method [^1] While we use the two-phase multi-task training pipeline (sequential) in our main experiments, in Appendix B, we show that an equivalent single-phase multi-task training pipeline (joint) results in similar performance. for UDA with masked language modeling. In the first phase, we take a pre-trained auto-regressive language model and perform unsupervised training through masked language modeling, on the combination of unlabeled data from both the source and target domains. In the second phase, we perform supervised fine-tuning on the downstream classification task using labeled data from the source domain.
Task 1: MLM Pre-training
@@
We aim to utilize a pre-trained autoregressive language model for continuing pre-training on unlabeled data from both the source and target domains, with the masked language modeling task. Here, we reuse all the input sequences from the source-labeled dataset $\mathcal{D}_{\text{src}}$ as the source-unlabeled dataset, denoted as $\mathcal{D}_{\text{src}}^{x}$. Next, similar to Raffel et al. (2020), for any unlabeled sequence $x\in\mathcal{D}_{\text{src}}^{x}$ and $\mathcal{D}_{\text{src}}$, we use a prompt template to convert the sequence $x$ to an input-output sequence pair, i.e., $\mathbb{M}(x)=(\tilde{x},\tilde{y})$. Here, the prompt template first randomly masks words in the input sequence $x$ and prepends an instruction (i.e., “Fill in the blanks:”) to create a new input sequence $\tilde{x}$. An output sequence $\tilde{y}$ is then constructed by concatenating all the masked words separated by a special <sep> token.
For example,
[…]
$\tilde{y}=\text{{<sep>} movie was so {<sep>} of fun. {<sep>}}$>
@@
Given a pair $(\tilde{x},\tilde{y})$, we use an autoregressive LM parameterized by $\theta$ to compute the negative log-likelihood loss averaged over output words:
$\ell(\tilde{x},\tilde{y};\theta)=-\frac{1}{|\tilde{y}|}\sum_{t}\log P_{\theta}(\tilde{y}_{t}|\tilde{x},\tilde{y}_{1:t-1}).$ (1)
Notably, we convert the MLM task into a next-token prediction task by instructing an autoregressive LM to predict the output words, which allows us to reuse all the parameters of the LM without adding any new randomly-initialized parameters. Finally, we define the total loss on the combined unlabeled dataset $\mathcal{D}=\mathcal{D}_{\text{src}}^{x}\cup\mathcal{D}_{\text{tgt}}$ as:
$\mathcal{L}_{\text{MLM}}(\mathcal{D};\theta)={\frac{1}{|\mathcal{D}|}}\sum_{x\in\mathcal{D}}\ell(\mathbb{M}(x);\theta).$ (2)
While the above formulation uses the MLM task for continued pre-training, the use of an autoregressive LM allows for an easy extension to the Causal Language Modeling (CLM) task, which we demonstrate in Section 6.3.
@@
In the second phase, we only use labeled data from the source domain to fine-tune the model for the downstream classification task. Similar to the first phase, we use a prompt template [^2] Prompt templates were selected from the Public Pool of Prompts Bach et al. (2022). to create input-output sequence pairs from the labeled data in the source domain. Specifically, for a labeled example $(x,y)\in\mathcal{D}_{\text{src}}$, a new prompt template appends to $x$ an instruction for prompting the LM to perform classification, and converts the label $y$ to its corresponding text description $\tilde{y}$, i.e., $\mathbb{C}(x,y)=(\tilde{x},\tilde{y})$.
For example, a labeled example with the positive sentiment from a sentiment classification task is converted as follows:
$x=\text{I like this movie.}~{}~{}~{}~{}~{}y=1$
$\tilde{x}=[x]\text{ Is this sentence positive or negative?}$?
$\tilde{y}=\text{Positive}$
Given the augmented sequence pair $(\tilde{x},\tilde{y})$ and the model trained after the first phase, we compute the same negative log-likelihood loss $\ell(\tilde{x},\tilde{y};\theta)$ in Eq. (1). This again allows us to reuse all model parameters, including the language model prediction head. Finally, we define the total loss on the source-labeled dataset in the second phase as:
@@
Parameter-Efficient Fine-Tuning (PEFT)
The above formulation is general and can be applied to fine-tune all model parameters. Additionally, our FEUDA framework is compatible with the parameter-efficient fine-tuning approach. The PEFT approach is desirable because it adds only a small amount of learnable parameters $\phi$ to a pre-trained language model $\theta$, and fine-tunes only $\phi$ to perform prediction while keeping the other model parameters $\theta$ frozen. We utilize two instantiations in our implementations: Adapters Houlsby et al. (2019) and (IA)${}^{3}$ Liu et al. (2022). Both are high-performing PEFT approaches, with (IA)${}^{3}$ using fewer learnable parameters. More details about both methods can be found in Appendix A.
5 Experimental Setup
5.1 Datasets
We follow the setup from Malik et al. (2023), and use two sentence classification datasets with 5 domains each. This results in a total of 40 pairs of source and target domains. Appendix C shows more details about the evaluation benchmarks.
Source normal-$\rightarrow$ Target Source Accuracy Target Accuracy
Src Only Src+Tgt MMD FEUDA Src Only Src+Tgt MMD FEUDA
T->F 78.4 (2.2) 79.1 (0.5) 78.3 (0.1) 80.5 (0.1) 73.7 (1.1) 77.2 (0.4) 69.7 (0.8) 74.1 (0.9)
T->G 78.4 (2.2) 79.8 (0.1) 78.2 (0.2) 80.1 (0.3) 73.7 (1.1) 83.6 (0.7) 79.3 (0.5) 83.6 (0.3)
T->S 78.4 (2.2) 80.6 (3.5) 79.8 (0.2) 79.8 (0.4) 73.7 (1.1) 72.3 (0.5) 69.6 (0.1) 70.7 (0.6)
T->Te 78.4 (2.2) 79.2 (0.4) 78.0 (0.0) 81.1 (0.0) 74.5 (2.2) 77.8 (0.1) 69.4 (0.8) 76.8 (0.0)
F->T 76.0 (0.2) 77.6 (0.3) 72.9 (0.2) 67.6 (1.5) 75.6 (0.7) 79.9 (0.1) 69.9 (0.2) 65.4 (1.8)
F->G 76.0 (0.2) 77.6 (0.6) 53.3 (21.4) 73.2 (2.3) 75.6 (0.7) 82.3 (0.1) 54.3 (23.4) 78.8 (2.5)
F->S 76.0 (0.2) 77.4 (0.3) 69.7 (2.7) 69.8 (1.8) 75.6 (0.7) 72.1 (0.2) 64.6 (1.8) 65.3 (1.6)
F->Te 76.0 (0.2) 77.8 (0.6) 70.6 (0.9) 74.4 (0.4) 75.6 (0.7) 78.3 (0.6) 64.6 (0.7) 72.5 (0.2)
G->T 82.1 (0.3) 83.6 (0.4) 80.9 (0.7) 82.3 (0.8) 73.0 (0.0) 79.9 (0.4) 75.9 (0.3) 75.8 (0.6)
G->F 82.1 (0.3) 81.6 (0.1) 79.8 (0.7) 81.7 (0.2) 73.0 (0.0) 76.7 (0.1) 69.9 (0.2) 73.5 (0.2)
G->S 82.1 (0.3) 82.9 (0.1) 80.9 (0.0) 79.8 (1.7) 73.0 (0.0) 73.1 (0.0) 69.4 (0.1) 68.0 (1.8)
G->Te 82.1 (0.3) 83.2 (0.2) 80.1 (0.1) 82.1 (0.1) 73.0 (0.0) 78.1 (0.6) 69.9 (0.3) 73.5 (0.6)
S->T 70.9 (1.7) 71.9 (0.1) 69.1 (0.9) 71.2 (0.2) 72.9 (1.5) 79.5 (0.3) 74.4 (1.7) 76.8 (1.0)
S->F 70.9 (1.7) 71.6 (0.4) 70.4 (0.0) 70.3 (0.7) 72.9 (1.5) 77.7 (0.2) 73.1 (0.0) 72.4 (0.7)
S->G 70.9 (1.7) 72.8 (0.2) 68.5 (0.8) 66.4 (1.5) 72.9 (1.5) 83.4 (0.2) 78.2 (0.5) 76.3 (0.9)
S->Te 70.9 (1.7) 73.3 (0.0) 67.5 (1.4) 71.0 (1.7) 72.9 (1.5) 78.5 (0.0) 66.7 (0.2) 74.8 (1.3)
Te->T 77.5 (0.2) 78.2 (0.4) 75.5 (0.5) 78.7 (0.2) 74.9 (0.2) 79.8 (0.3) 71.4 (0.0) 76.5 (0.4)
Te->F 77.5 (0.2) 78.1 (0.4) 75.2 (0.7) 77.1 (0.0) 74.9 (0.2) 77.9 (0.1) 69.9 (0.5) 74.3 (0.5)
Te->G 77.5 (0.2) 78.6 (0.2) 74.8 (0.5) 78.8 (0.1) 74.9 (0.2) 82.5 (0.1) 75.6 (1.6) 82.0 (0.6)
Te->S 77.5 (0.2) 78.7 (0.0) 75.3 (0.1) 78.8 (0.4) 74.9 (0.2) 72.2 (0.0) 68.0 (0.4) 71.3 (0.5)
Table 1: Comparison of FEUDA and MMD by classification accuracy on the MNLI dataset, using the T5v1.1 base model, and (IA)${}^{3}$ PEFT method. FEUDA is competitive with MMD, often outperforming it. The highest values between FEUDA and MMD have been marked in bold. Cases where Src Only outperforms both FEUDA and MMD on the target have been underlined. However, it must be noted that in a majority of these cases the upper bound Src+Tgt is comparable to or weaker than Src Only, indicating noise in the domain pair.
MNLI
The MNLI corpus Williams et al. (2018) contains sentence pairs across multiple genres: Travel (T), Fiction (F), Government (G), Slate (S), and Telephone (Te). The task classifies every sentence pair as entailment, neutral, or contradiction.
Amazon
The Multi-Domain Sentiment Analysis Dataset Blitzer et al. (2007) contains Amazon product reviews for different types of products. We use reviews from the Apparel (A), Baby (B), Books (Bo), Cameras (C), and Movies (M) domains. Each review is labeled as either positive or negative.
5.2 Models and Training
Models
For our main experiments, we use T5v1.1, which is an improved version of the original T5 model Raffel et al. (2020). Unlike the original T5 model, T5v1.1 is not trained on any supervised datasets. We use the base (60M parameters) and XL (3B parameters) versions of the model. We also use T0 (3B parameters), which has been optimized for zero-shot generalization by training on supervised natural language prompts Sanh et al. (2022). Furthermore, to test the sensitivity of FEUDA across architectures and masking styles, we also use GPT-2 medium (345M parameters) Radford et al. (2019), from the class of autoregressive decoder-only language models.
Source normal-$\rightarrow$ Target Source Accuracy Target Accuracy
Src Only Src+Tgt MMD FEUDA Src Only Src+Tgt MMD FEUDA
A->B 93.7 (0.1) 93.8 (0.3) 94.3 (0.2) 93.1 (0.3) 93.3 (0.4) 94.7 (0.2) 93.8 (0.3) 93.9 (0.3)
A->B 93.7 (0.1) 94.2 (0.1) 93.8 (0.1) 92.8 (0.6) 90.8 (0.6) 94.3 (0.4) 92.5 (1.1) 90.2 (1.2)
A->C 93.7 (0.1) 93.4 (0.3) 95.0 (0.0) 93.9 (0.5) 91.9 (0.1) 95.0 (0.2) 91.8 (0.5) 92.1 (0.5)
A->M 93.7 (0.1) 94.1 (0.3) 94.7 (0.3) 93.5 (0.4) 81.3 (1.4) 85.8 (0.5) 81.3 (0.6) 83.3 (0.5)
Table 2: Comparison of FEUDA and MMD by classification accuracy on the Amazon Product Review dataset, using the T5v1.1 base model, and (IA)${}^{3}$ PEFT method. FEUDA is competitive with MMD, often outperforming it. The highest values between FEUDA and MMD on the target domain have been marked in bold.

### v2

# How Useful is Continued Pre-Training for Generative Unsupervised Domain Adaptation?

@@ Abstract
Recent breakthroughs in scale have enabled the emergence of powerful generative language models, and the ability to fine-tune these models on various tasks by casting them into prompts or instructions. In this landscape, the problem of Unsupervised Domain Adaptation (UDA), or the problem of leveraging knowledge from a labeled source domain to an unlabeled target domain, has been left behind, with recent UDA methods still addressing discriminative classification. In particular, two popular UDA approaches, involving Continued Pre-Training (CPT) and learning domain invariant representations, have been under-explored in the generative setting, signaling a gap. In this work, we evaluate the utility of CPT for generative UDA. We first perform an empirical evaluation to measure the trade-offs between CPT and strong methods promoting domain invariance. We further evaluate how well the benefits of CPT extend to different architectures, tuning methods and data regimes. We then motivate the use of CPT by studying to what degree it benefits classification performance on the target domain. Finally, we attempt to understand the mechanism behind which CPT improves classification performance on the unlabeled target domain. Our findings suggest that a implicitly learns the downstream task while predicting masked words informative to that task. Our work connects the body of UDA research with that of instruction tuning, enabling an initial step towards a wider applicability of modern language models.
@@ 1 Introduction
Recent advancements in the pre-training of language models have enabled the widespread use of powerful generative models, which can be leveraged across multiple domains with no training (Brown et al., 2020; Scao et al., 2022; Touvron et al., 2023, inter alia). Despite these advancements, these autoregressive models are still fragile under certain kinds of data distribution shifts, making their applications across these domains challenging (Ribeiro et al., 2020). This is addressed, in part, by the concept of instruction tuning with templates (Zhang et al., 2023; Sanh et al., 2022; Ouyang et al., 2022; Wang et al., 2022; Wei et al., 2022), enabling the learning of new tasks without any randomly initialized parameters.
The problem of unsupervised domain adaptation (UDA) leverages learned knowledge from a labeled source domain to an unlabeled target domain (Pan & Yang, 2010; Ganin & Lempitsky, 2015; Long et al., 2015, inter alia). It is useful for adaptation to unlabeled domains with high labeling costs, where supervised instruction tuning does not suffice. Despite the pervasive need for models to generalize to such domains, recent UDA methods still address discriminative classification, barring the application of these approaches to recent generative models. In particular, Continued Pre-Training and Domain Invariance based methods, two widely popular classes of UDA approaches (Ramponi & Plank, 2020), are completely unexplored in the generative setting.
The Continued Pre-Training (CPT) approach involves extended pre-training on a domain or task, followed by supervised training on the downstream task (Gururangan et al., 2020). This approach has been widely used for adaptation to labeled domains (Gao et al., 2021; Kim et al., 2021; Hung et al., 2023, inter alia), and in the UDA setup for unlabeled domains (Han & Eisenstein, 2019; Zhang et al., 2021b; Karouzos et al., 2021; Pfeiffer et al., 2020; Parović et al., 2023). Invariance based approaches attempt to learn representations that are invariant across domains (Tzeng et al., 2014; Ganin et al., 2016; Wu & Shi, 2022; Guo et al., 2022), with the notion that when the learned representations from both domains cannot be distinguished by a classifier and the classifier performs well on the source domain, it will also exhibit strong performance on the target domain. These two classes of methods introduce a trade-off: invariance based methods suffer from instability issues (Han & Eisenstein, 2019; Kashyap et al., 2021), while continued pre-training requires a larger computational budget. But how would this trade-off extrapolate to the generative setting? For example, invariance based methods are well motivated in discriminative tasks, where there is a clear decision boundary; however, the same does not hold for generative tasks.
To address these gaps, we introduce the setting of Generative UDA, where an autoregressive model is trained to leverage knowledge from a labeled source domain to an unlabeled target domain, using only next word prediction as its objective. We formalize the use of CPT for this setting in Section 2, and then attempt to explore and understand the behaviour of CPT for Generative UDA. We begin by performing an empirical analysis on 40 real-world domain pairs to explore the tradeoff between continued pre-training and invariance based approaches, and find vanilla CPT to be competitive with and significantly more stable than a state of the art invariance based approach (Section 3). We then stress test CPT, by applying it across varying model architectures and scales, tuning approaches and data regimes. We find that CPT is robust across these settings, unlike our invariance based approach (Section 4).
With recent language models being trained across vast corpora which may include domains similar to the target domain, the requirement for continued pre-training may be raised to question. In Section 5, we show that continued pre-training is indeed essential for strong downstream performance on the target domain, and this performance rapidly degrades with limited target domain exposure. Finally, we attempt to shed light on how masking plays a role in improving classification accuracy on the unlabeled target domain in Section 6. We find that the model may implicitly learn the downstream task as it predicts masked words that are informative to the downstream task.
Our work attempts to connect the body of UDA research with recent trends in language modeling, by providing a set of insights into the behaviour of the popular class of continued pre-training approaches, in the Generative UDA setting. We hope this enables an initial step towards a wider applicability of modern language models.
2 Continued Pre-Training for Generative UDA
2.1 Preliminaries: The UDA Problem
@@
We consider a text classification task, where $\mathcal{X}$ is the input space of all text sentences and $\mathcal{Y}=\{1,...K\}$ is the label space. In the UDA problem, we have access to a source labeled dataset $\mathcal{D}_{\text{src}}=\{(x_{i},y_{i})\}_{i=1}^{N}$ consisting of samples from a joint distribution $P_{\text{src}}$, and a target unlabeled dataset $\mathcal{D}_{\text{tgt}}=\{x_{j}\}_{j=1}^{M}$ sampling from a target input distribution $P^{\mathcal{X}}_{\text{tgt}}$. We further denote $P^{\mathcal{X}}_{\text{src}}$ as the marginal distribution of $P_{\text{src}}$ on $\mathcal{X}$, where $P^{\mathcal{X}}_{\text{src}}\neq P^{\mathcal{X}}_{\text{tgt}}$. The goal of UDA is to learn a function $f:\mathcal{X}\to\mathcal{Y}$ that minimizes the error rate $\mathbb{E}_{x\sim P^{\mathcal{X}}_{\text{tgt}}}\;\mathbbm{1}[f(x)\neq y]$.
2.2 CPT for UDA as a Sequence of Prompt Based Tasks
@@
We now formalize the extension of continued pre-training to the setting of generative UDA. We use the traditional two-phase training pipeline from Gururangan et al. (2020) [^1] While we use the two-phase multi-task training pipeline (sequential) in our main experiments, in Appendix I, we show that an equivalent single-phase multi-task training pipeline (joint) results in similar performance.. The first phase uses templates to cast the source and target domain sequences into an autoregressive pre-training task. [^2] We investigate mask language modeling for T5 models and switch to causal language modeling for decoder-only models with a few simple template changes. We compare both in Section 4. The second phase applies supervised instruction-tuning of the model on source-labeled data.
Task 1: Autoregressive Continued Pre-training
@@
We reuse the input sequences from the source-labeled dataset $\mathcal{D}_{\text{src}}$ as the source-unlabeled dataset, denoted as $\mathcal{D}_{\text{src}}^{\mathcal{X}}$. Next, similar to Raffel et al. (2020); Song et al. (2019), for an unlabeled sequence $x\in\mathcal{D}_{\text{src}}^{\mathcal{X}}$ and $\mathcal{D}_{\text{src}}$, we use a prompt template to convert the sequence $x$ to an input-output sequence pair, i.e., $\mathbb{M}(x)=(\tilde{x},\tilde{y})$. For masked language modeling (MLM), an instruction is pre-pended to a randomly masked sequence $x$ to create $\tilde{x}$. The output sequence $\tilde{y}$ is a concatenation of masked words from $x$. For example, given $x=$“The movie was so cool! Two hours of fun.”, we construct $\tilde{x}=$“Fill in the blanks: ”The _ cool! Two hours _”, and $\tilde{y}=$ <sep> movie was so <sep> of fun. <sep>. For causal language modeling (CLM), $\tilde{x}=x$.
@@
Given $(\tilde{x},\tilde{y})$, we train an autoregressive LM parameterized by $\theta$ to minimize the negative log-likelihood loss averaged over output words and the total loss over a corpus $\mathcal{D}=\mathcal{D}_{\text{src}}^{\mathcal{X}}\cup\mathcal{D}_{\text{tgt}}$.
$\ell(\tilde{x},\tilde{y};\theta)=-\frac{1}{|\tilde{y}|}\sum_{t}\log P_{\theta}(\tilde{y}_{t}|\tilde{x},\tilde{y}_{1:t-1});~{}~{}~{}~{}~{}~{}\mathcal{L}_{\text{CPT}}(\mathcal{D};\theta)={\frac{1}{|\mathcal{D}|}}\sum_{x\in\mathcal{D}}\ell(\mathbb{M}(x);\theta).$ (1)
@@
In the second phase, we use labeled data from the source domain to train the model on the downstream classification task. Similar to the first phase, we use prompts [^3] Prompt templates were selected from the Public Pool of Prompts (Bach et al., 2022). to generate input-output sequence pairs: $\mathbb{C}(x,y)=(\tilde{x},\tilde{y})\;\forall\;(x,y)\in\mathcal{D}_{\text{src}}$. For example, for sentiment classification, if $x=$“I like this movie.”, $y=1\Rightarrow$ $\tilde{x}=$ “[$x$] Is this sentence positive or negative?”, $\tilde{y}=$ “Positive”.
Given the augmented sequence pair $(\tilde{x},\tilde{y})$ and the model trained after the first phase, we compute the same negative log-likelihood loss $\ell(\tilde{x},\tilde{y};\theta)$ in Eq. (1). Finally, we define the total loss on the source-labeled dataset in the second phase as:
@@
3 Evaluating the Efficacy of Continued Pre-training for Generative UDA
3.1 Experimental Setup
Datasets
We use the MNLI and Amazon Review classification datasets, which are widely used UDA benchmarks (Malik et al., 2023; Karouzos et al., 2021; Guo et al., 2020). The MNLI corpus (Williams et al., 2018) contains sentence pairs across five genres: Travel (T), Fiction (F), Government (G), Slate (S), and Telephone (Te). The task classifies every sentence pair as entailment, neutral, or contradiction. The Multi-Domain Sentiment Analysis Dataset (Blitzer et al., 2007) contains binary sentiment reviews for different types of Amazon products. We use reviews from the Apparel (A), Baby (B), Books (Bo), Cameras (C), and Movies (M) domains. We evaluate a total of 40 pairs of source and target domains, across the two datasets. Appendix A contains more details about the datasets.
Models and Tuning Methods
Our main experiments use the T5v1.1 base model and (IA)3 (Liu et al., 2022) PEFT method. T5v1.1 is an improved version of the original T5 model (Raffel et al., 2020), and unlike the original T5 model, it is not trained on any supervised datasets. We then extend our evaluation to different model architectures (T0, GPT-2), tuning methods (full fine-tuning, adapters) and data regimes (Section 4).